OpenStudy (twalt13):

Help solve please. Write the sum using summation notation, assuming the suggested pattern continues.1 - 3 + 9 - 27 + ...

OpenStudy (anonymous):

i think this is correct

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} (-3)^{n}\]

OpenStudy (twalt13):

\[\sum_{n=0}^{\infty}1(-3)^{n}\]

OpenStudy (anonymous):

that 1 is irrelevant because \[1*n=n\]

OpenStudy (twalt13):

okay thanks